OpenStudy (anonymous):

Use the Pythagorean theorem to find one leg if the hypotenuse is 13 and the other leg is 5.

OpenStudy (anonymous):

12 3 13.93 11.24 ?

OpenStudy (anonymous):

hyp^2 = side^2 + side^2 13^2 = 5^2 + side^2 169 = 25 + side^2 144 = side ^2 12 = side

OpenStudy (anonymous):

Thank you so much!

OpenStudy (anonymous):

this is one of those right-angled triangles whose sides are whole numbers. the lowest is a triangle with sides 3,4 and 5.